# A new perspective on modulus stabilization

### v2

@@ V.1 Global minimum for $V_{b}(\Phi)$
The existence of a global minimum for the bulk scalar potential has strong implications for the OR solution of $\Phi$ (see Eq.23). Particularly, the OR equation of Eq.(23) indicates the existence of a stable fixed point for $\Phi_{OR}$ at $\Phi_{0}$ the scalar field value at which $V_{b}(\Phi)$ attains its global minimum. As a result, the profile of $\Phi_{OR}$ is such that it gets a lot of contribution from the global minimum well, around $\Phi_{0}$. This allows us to treat the bulk potential as an expansion around its minimum. As we are interested in the nature of the potential, we can take the global minimum at the origin i.e. $\Phi_{0}=0$ and $V(\Phi_{0})=0$. The rescaling essentially implies corrections to the bulk cosmological constant which is assumed to still remain negative ($AdS_{5}$). We alternatively call $V_{b}(\Phi)$ as $V(\Phi)$ and $V_{b}^{\prime\prime}(\Phi)|_{\Phi=\Phi_{0}}$ as $V^{\prime\prime}$ for brevity.
@@
V.1.1 GW conditions
With GW conditions [as defined in Sec.II(C)], the analysis is exactly the same as Sec.II(C) except $V^{\prime\prime}(\Phi)|_{\Phi=\Phi_{0}}$ representing the squared mass of the bulk scalar. Hence, we get modulus stabilization.
Under GW conditions, if the bulk scalar potential has a global minimum satisfying $V^{\prime\prime}>0$, then it implies modulus stabilization.
V.1.2 CMS conditions
@@
$\begin{split}V(r_{c})=\frac{k^{4}\alpha^{2}}{4}e^{-8kr_{c}\pi}(e^{4kr_{c}\pi}-1)\\ -\frac{(V^{\prime\prime})^{2}}{8k}\frac{k^{3}v^{2}}{V^{\prime\prime}+8k^{2}}(e^{-(\frac{r_{c}V^{\prime\prime}}{2k}+4kr_{c})\pi}-1)\\ -k^{4}\alpha ve^{-4kr_{c}\pi}(e^{-r_{c}V^{\prime\prime}\pi/4k}-1)\\ -2k^{4}v^{2}\frac{V^{\prime\prime}}{V^{\prime\prime}+8k^{2}}(e^{-(\frac{r_{c}V^{\prime\prime}}{2k}+4kr_{c})\pi}-1)\\ -\frac{k^{4}\alpha^{2}}{2}e^{-4kr_{c}\pi}+2k^{4}\alpha ve^{-(\frac{r_{c}V^{\prime\prime}}{4k}+4kr_{c})\pi}\end{split}$ (64)
@@
$\begin{split}V^{\prime}(r_{c})=k^{5}\alpha^{2}\pi e^{-4kr_{c}\pi}-4k^{5}\alpha\pi ve^{-4kr_{c}\pi}+2k^{5}\alpha^{2}\pi e^{-8kr_{c}\pi}\\ +\frac{k\pi}{16}v^{2}(V^{\prime\prime})^{2}e^{-(\frac{r_{c}V^{\prime\prime}}{2k}+4kr_{c})\pi}+k^{4}\alpha\pi v(4k+\frac{V^{\prime\prime}}{4k})e^{-(\frac{r_{c}V^{\prime\prime}}{4k}+4kr_{c})\pi}\\ +k^{3}v^{2}\pi V^{\prime\prime}e^{-(\frac{r_{c}V^{\prime\prime}}{2k}+4kr_{c})\pi}-2\alpha k^{4}v\pi(\frac{V^{\prime\prime}}{4k}+4k)e^{-(\frac{r_{c}V^{\prime\prime}}{4k}+4kr_{c})\pi}\end{split}$ (65)
The fact that $14kr_{c}>>$ 1 allows us to neglect the $e^{-8kr_{c}\pi}$ term. At the extremum (supposing it exists) say $r_{c}^{*}$, Eq.(65) vanishes. The second derivative of the radion potential at extremum $r_{c}^{*}$ is then given by
$\begin{split}V^{\prime\prime}(r_{c})|_{r_{c}^{*}}=\frac{k^{2}\pi^{2}\alpha v}{16}V^{\prime\prime}(V^{\prime\prime}+16k^{2})e^{-(\frac{r_{c}V^{\prime\prime}}{4k}+4kr_{c})\pi}\end{split}$ (66)
@@
V.2 Global maximum but no global minimum for $V_{b}(\Phi)$
@@
The analysis follows as same except now, the perturbative solution for the scalar field is a runaway solution as the point of maximum is an unstable fixed point of the system and there exists no global minimum. The bulk potential can be approximated as same except $V^{\prime\prime}\leq 0$ for the proof, we take it to be $V^{\prime\prime}<0$.
V.2.1 GW conditions
The analysis is as given in Sec.IV(A). We deal with a negative mass term which essentially results in modulus destabilization as we showed previously.
@@ V.2.1 GW conditions
Under GW conditions, if the bulk scalar potential has a global maximum satisfying $V^{\prime\prime}>0$, then the corresponding radion potential has no extremum for a positive value of $r_{c}$ implying no modulus stabilization.
V.2.2 CMS conditions
@@
$V^{\prime\prime}>0$ allows us to only consider the terms involving $e^{-\frac{r_{c}V^{\prime\prime}\pi}{4k}}$ and $e^{-\frac{r_{c}V^{\prime\prime}\pi}{2k}}$ in Eq.(65) as they are increasing exponentials, $V^{\prime\prime}$ being negative. Supposing the extremum exists, then the second derivative of the radion potential at the extremum $r_{c}^{*}$ is given by
$V^{\prime\prime}(r_{c})|_{r_{c}^{*}}=-\frac{k^{2}\pi^{2}\alpha v}{8}V^{\prime\prime}(V^{\prime\prime}+16k^{2})e^{-(\frac{r_{c}V^{\prime\prime}}{4k}+4kr_{c})\pi}$ (67)
@@
where we have used the vanishing of the first derivative at extremum. $V^{\prime\prime}$ being negative and ensuring $\alpha v>0$ as in the previous proof, we have two cases: (a) $V^{\prime\prime}<-16k^{2}$ indicates a maximum and (b) $-16k^{2}<V^{\prime\prime}<0$ indicates a minimum. Thus we get to the following theorem,
@@
Under CMS conditions, if the bulk scalar potential has a global maximum but no global minimum, and the corresponding radion potential has an extremum for a positive value of $r_{c}$, then the extremum is a maximum of the radion potential if $V^{\prime\prime}<-16k^{2}$ or a minimum of the radion potential if $-16k^{2}<V^{\prime\prime}<0$ implying modulus destabilization or stabilization respectively.
Though the $V^{\prime\prime}=0$ case was not included here, we suspect that the above results hold for such scenarios as well. The case study of Sec.IV(C) supports this belief.
Thus, from these, we can conclude that if the bulk GW potential has a global minimum, then modulus stabilization is guaranteed. However, for GW bulk potentials with a global maximum and no global minimum, we saw that modulus stabilization may or may not be achieved depending on the choice of our parameters for the bulk potential. Note that these deductions are consistent with but do not explain the case studies for bulk cubic interaction and the Mishra-Randall ansatz.
It must be noted that we went for a full expansion of the radion potential for the CMS proofs unlike Ref.[5] which calculates only to the leading order in $v$ albeit incorporating gravitational sector contribution.
@@
VI Criteria for bulk potential to ensure modulus stabilization
@@
In this section, we will frequently use tools from Sec. III. Firstly, note that for a general bulk scalar potential $V(\Phi)$ and defining $z=r_{c}\phi$, the general scalar solution is given by
@@
We write $V^{\prime}(\Phi)|_{\pi r_{c}}$ as $V^{\prime}(\Phi)$ and $V^{\prime}(\Phi_{OR})|_{\pi r_{c}}$ as $V^{\prime}(\Phi_{OR})$ for brevity. Then, plugging these along with the brane potentials defined in Sec.III into Eq.(19), applying Leibniz’s rule to get the first derivative of the radion potential $U(r_{c})$ and setting it to zero (we demand the existence of an extremum at $r_{c}^{*}$) gives us the following relation evaluated at $r_{c}^{*}$
$\begin{split}\frac{\pi}{2}k^{5}\alpha^{2}+\frac{\pi(V^{\prime}(\Phi_{OR}))^{2}}{32k^{2}r_{c}^{*2}}+\frac{\pi k^{3/2}\alpha V^{\prime}(\Phi_{OR})}{4r_{c}^{*}}-\pi k^{5/2}\alpha V^{\prime}(\Phi)\\ -\frac{\pi V^{\prime}(\Phi)V^{\prime}(\Phi_{OR})}{4kr_{c}^{*}}+2\pi\alpha^{2}k^{5}-8\alpha\pi k^{7/2}\Phi_{OR}(\pi r_{c}^{*})\\ +2\alpha k^{5/2}\frac{d\Phi_{OR}(\pi r_{c})}{dr_{c}}|_{r_{c}^{*}}=0\end{split}$ (69)
@@
We now evaluate $U^{\prime\prime}(r_{c})$ at $r_{c}^{*}$.
$\begin{split}U^{\prime\prime}(r_{c})|_{r_{c}^{*}}=e^{-4kr_{c}^{*}\pi}[\frac{V^{\prime\prime}(\Phi_{OR})}{4r_{c}^{*}}.\frac{d\Phi_{OR}(\pi r_{c})}{dr_{c}}\\ -kV^{\prime\prime}(\Phi)\frac{d\Phi(\pi r_{c})}{dr_{c}}-8k^{2}\frac{d\Phi_{OR}(\pi r_{c})}{dr_{c}}\\ +\frac{2k}{\pi}\frac{d^{2}\Phi_{OR}(\pi r_{c})}{dr_{c}^{2}}-\frac{1}{4r_{c}^{*2}}V^{\prime}(\Phi_{OR})]\end{split}$ (71)
@@
$\begin{split}\frac{2k}{\pi}\frac{d^{2}\Phi_{OR}(\pi r_{c})}{dr_{c}^{2}}|_{r_{c}^{*}}>[8k^{2}+kV^{\prime\prime}(\Phi)|_{r_{c}^{*}}\\ -\frac{V^{\prime\prime}(\Phi_{OR})|_{r_{c}^{*}}}{4r_{c}^{*}}]\frac{d\Phi_{OR}(\pi r_{c})}{dr_{c}}|_{r_{c}^{*}}+\frac{1}{4r_{c}^{*2}}V^{\prime}(\Phi_{OR})|_{r_{c}^{*}}\end{split}$ (72)
@@
$\begin{split}(\frac{2k}{\pi})^{2}\frac{d^{2}\Phi_{OR}(\pi r_{c})}{dr_{c}^{2}}|_{r_{c}^{*}}>[8k^{2}+kV^{\prime\prime}(\Phi)-\frac{V^{\prime\prime}(\Phi_{OR})}{4r_{c}^{*}}]\\ [kV^{\prime}(\Phi)+8k^{2}\Phi_{OR}(\pi r_{c}^{*})-\frac{V^{\prime}(\Phi_{OR})}{4r_{c}^{*}}-\frac{5}{2}\alpha k^{7/2}]\\ +\frac{k}{2\pi r_{c}^{*2}}V^{\prime}(\Phi_{OR})\end{split}$ (73)
@@
where on the right hand side, the “evaluated at $r_{c}^{*}$ is implicit. As one can see, the Eqs.(70) and (73) are completely determined by the bulk scalar potential (note that $\Phi(z)$ depends only on $V^{\prime}(\Phi)$). Thus, these are the two relations that any general bulk potential must satisfy in order for the radion potential to have a stable minimum. It should again be noted that these results hold under our assumed boundary conditions.

### v4

@@ 5.1 Global minimum for $V_{b}(\Phi)$
The existence of a global minimum for the bulk scalar potential has strong implications for the OR solution of $\Phi$ (see Eq.3.1). Particularly, the OR equation of Eq.(3.1) indicates the existence of a stable fixed point for $\Phi_{OR}$ at $\Phi_{0}$ - the scalar field value at which $V_{b}(\Phi)$ attains its global minimum. As a result, the profile of $\Phi_{OR}$ is such that it gets a lot of contribution from the global minimum well, around $\Phi_{0}$. This allows us to treat the bulk potential as an expansion around its minimum. As we are interested in the nature of the potential, we can take the global minimum at the origin i.e. $\Phi_{0}=0$ and $V(\Phi_{0})=0$. The rescaling essentially implies corrections to the bulk cosmological constant which is assumed to still remain negative ($AdS_{5}$). We alternatively call $V_{b}(\Phi)$ as $V(\Phi)$ and $V_{b}^{\prime\prime}(\Phi)|_{\Phi=\Phi_{0}}$ as $V^{\prime\prime}$ for brevity.
@@
We note that the GW scalar field $\Phi$ in five dimensions has mass dimension $\Phi]=3/2$. As a result, higher-order terms in the bulk potential, such as $\Phi^{n}$ for $n\geq 4$ are suppressed by powers of the cutoff scale - the 5D Planck scale $M$. For instance, the operator $\Phi^{4}$ has mass dimension 6 and thus appears with a coefficient of order $1/M$, while higher order terms are even more suppressed. Therefore, in the effective field theory regime where $\Phi<<M^{3/2}$, the potential can reliably be truncated at quadratic (or at most cubic) order.
With GW conditions [as defined in Sec.2.3], the analysis is exactly the same as Sec.2.3 except $V^{\prime\prime}(\Phi)|_{\Phi=\Phi_{0}}$ representing the squared mass of the bulk scalar. Hence, we can say- under GW conditions, if the bulk scalar potential has a global minimum satisfying $V^{\prime\prime}>0$, then it implies modulus stabilization. However, an extensive calculation with the CMS conditions is yet to be worked out which is what we do now. It must be noted that we go for a full expansion of the radion potential unlike Ref.Chacko1 which calculates only to the leading order in $v$ albeit incorporating gravitational sector contribution.
@@
$\begin{split}V_{GW}(r_{c})=\frac{k^{4}\alpha^{2}}{4}e^{-8kr_{c}\pi}(e^{4kr_{c}\pi}-1)-\frac{(V^{\prime\prime})^{2}}{8k}\frac{k^{3}v^{2}}{V^{\prime\prime}+8k^{2}}(e^{-(\frac{r_{c}V^{\prime\prime}}{2k}+4kr_{c})\pi}-1)\\ -k^{4}\alpha ve^{-4kr_{c}\pi}(e^{-r_{c}V^{\prime\prime}\pi/4k}-1)-2k^{4}v^{2}\frac{V^{\prime\prime}}{V^{\prime\prime}+8k^{2}}(e^{-(\frac{r_{c}V^{\prime\prime}}{2k}+4kr_{c})\pi}-1)\\ -\frac{k^{4}\alpha^{2}}{2}e^{-4kr_{c}\pi}+2k^{4}\alpha ve^{-(\frac{r_{c}V^{\prime\prime}}{4k}+4kr_{c})\pi}\end{split}$ (60)
@@
$\begin{split}V^{\prime}_{GW}(r_{c})=k^{5}\alpha^{2}\pi e^{-4kr_{c}\pi}-4k^{5}\alpha\pi ve^{-4kr_{c}\pi}+2k^{5}\alpha^{2}\pi e^{-8kr_{c}\pi}\\ +\frac{k\pi}{16}v^{2}(V^{\prime\prime})^{2}e^{-(\frac{r_{c}V^{\prime\prime}}{2k}+4kr_{c})\pi}+k^{4}\alpha\pi v(4k+\frac{V^{\prime\prime}}{4k})e^{-(\frac{r_{c}V^{\prime\prime}}{4k}+4kr_{c})\pi}\\ +k^{3}v^{2}\pi V^{\prime\prime}e^{-(\frac{r_{c}V^{\prime\prime}}{2k}+4kr_{c})\pi}-2\alpha k^{4}v\pi(4k+\frac{V^{\prime\prime}}{4k})e^{-(\frac{r_{c}V^{\prime\prime}}{4k}+4kr_{c})\pi}\end{split}$ (61)
The fact that $4kr_{c}>>1$ allows us to neglect the $e^{-8kr_{c}\pi}$ term. If $V^{\prime\prime}\gtrsim 16k^{2}$, then one obtains $e^{-(\frac{r_{c}V^{\prime\prime}}{4k}+4kr_{c})\pi}\lesssim e^{-8kr_{c}\pi}$ and $e^{-(\frac{r_{c}V^{\prime\prime}}{2k}+4kr_{c})\pi}\lesssim e^{-12kr_{c}\pi}$. Consequently, it does not make sense, in general, to neglect the $e^{-8kr_{c}\pi}$ term alone, while simultaneously retaining the other terms. Thus, neglecting the $e^{-8kr_{c}\pi}$ term with respect to all the other terms in Eq. 5.6 is justified only for $V^{\prime\prime}<16k^{2}$. This upper bound for $V^{\prime\prime}$ has its justifications - firstly, $V^{\prime\prime}$ corresponds to mass squared of the bulk scalar and we take $\frac{V^{\prime\prime}}{k^{2}}<<1$, where k is the AdS curvature scale to ensure that backreaction can be neglected and we can solve for the bulk scalar in the original RS background (which is what we do here). Secondly, in the context of the AdS-Swampland conjecture lust2019ads, a consistent theory of quantum gravity is expected to forbid parametrically light or isolated AdS vacua. Specifically, the AdS curvature scale $k$ (with $\Lambda_{b}\sim-k^{2}$) is conjectured to be tied to the mass scale of an infinite tower of states, $m_{\text{tower}}\sim|\Lambda_{b}|^{1/2}\sim k$. Consequently, the effective field theory description breaks down when scalar masses in the bulk significantly exceed the AdS curvature scale, implying a soft upper bound $V^{\prime\prime}\lesssim\mathcal{O}(1)\,k^{2}$.
At the extremum (supposing it exists) say $r_{c}^{*}$, Eq.(61) vanishes. The expression for $r_{c}^{*}$ is given by,
$\begin{split}e^{-\frac{r_{c}^{*}V^{\prime\prime}\pi}{4k}}=\frac{k^{3}\alpha v(4k+\frac{V^{\prime\prime}}{4k})\pm\sqrt{X}}{V^{\prime\prime}(2kv^{2}+\frac{v^{2}V^{\prime\prime}}{8})}\end{split}$
where
$\begin{split}X=(V^{\prime\prime})^{2}(k^{4}\alpha v^{3}-\frac{3}{16}k^{4}\alpha^{2}v^{2})+V^{\prime\prime}(16k^{6}\alpha v^{3}-2k^{6}\alpha^{2}v^{2})+16k^{8}\alpha^{2}v^{2}\end{split}$
The constraint on the $\alpha,v,V^{\prime\prime}$ parameter space comes from two requirements: (a) $X\geq 0$ and (b) $0<e^{-\frac{r_{c}^{*}V^{\prime\prime}\pi}{4k}}<1$ to ensure the modulus value $r_{c}^{*}>0$ (as $V^{\prime\prime}>0$). For $V^{\prime\prime}<0$ (see Sec.5.2), the second criterion changes as $1<e^{-\frac{r_{c}^{*}V^{\prime\prime}\pi}{4k}}<\infty$. Thus, we can have either zero, one or two extrema. The second derivative of the radion potential at an extremum $r_{c}^{*}$ is then given by
$\begin{split}V^{\prime\prime}_{GW}(r_{c})|_{r_{c}^{*}}=\frac{k^{2}\pi^{2}\alpha v}{16}V^{\prime\prime}(V^{\prime\prime}+16k^{2})e^{-(\frac{r_{c}^{*}V^{\prime\prime}}{4k}+4kr_{c}^{*})\pi}\end{split}$ (62)
@@
5.2 Global maximum but no global minimum for $V_{b}(\Phi)$
@@
The analysis follows as same except now, the perturbative solution for the scalar field is a runaway solution as the point of maximum is an unstable fixed point of the system and there exists no global minimum. The bulk potential can be approximated as same except $V^{\prime\prime}\leq 0$ - for the proof, we take it to be $V^{\prime\prime}<0$.
@@
With GW conditions, the analysis is again as given in Sec.4.1. We deal with a negative mass term which essentially results in modulus destabilization as we showed previously. To put it in words; under GW conditions, if the bulk scalar potential has a global maximum satisfying $V^{\prime\prime}<0$, then the corresponding radion potential has no extremum for a positive value of $r_{c}$ implying no modulus stabilization.. We now move onto the CMS conditions.
@@
$V^{\prime\prime}<0$ allows us to only consider the terms involving $e^{-\frac{r_{c}V^{\prime\prime}\pi}{4k}}$ and $e^{-\frac{r_{c}V^{\prime\prime}\pi}{2k}}$ in Eq.(61) as they are increasing exponentials, $V^{\prime\prime}$ being negative. Supposing an extremum exists, then the second derivative of the radion potential at the extremum $r_{c}^{*}$ is given by
$V^{\prime\prime}_{GW}(r_{c})|_{r_{c}^{*}}=-\frac{k^{2}\pi^{2}\alpha v}{16}V^{\prime\prime}(V^{\prime\prime}+16k^{2})e^{-(\frac{r_{c}^{*}V^{\prime\prime}}{4k}+4kr_{c}^{*})\pi}$ (63)
@@
where we have used the vanishing of the first derivative at extremum. $V^{\prime\prime}$ being negative and ensuring $\alpha v>0$ as in the previous proof, we have two cases: (a) $V^{\prime\prime}<-16k^{2}$ indicates a maximum and (b) $16k^{2}<V^{\prime\prime}<0$ indicates a minimum. Thus we get to the following theorem,
@@
Under CMS conditions, if the bulk scalar potential has a global maximum but no global minimum satisfying $V^{\prime\prime}<0$, and the corresponding radion potential has an extremum for a positive value of $r_{c}$, then the extremum is a maximum of the radion potential if $V^{\prime\prime}<-16k^{2}$ or a minimum of the radion potential if $16k^{2}<V^{\prime\prime}<0$ implying modulus destabilization or stabilization respectively.
A point to note here is that the BF condition $\frac{m^{2}}{k^{2}}+4>0$ translates to $\frac{V^{\prime\prime}}{16k^{2}}>-\frac{1}{4}$. Thus if the BF condition were to be satisfied i.e. the bulk field were to be stable, then the $\frac{V^{\prime\prime}}{16k^{2}}<-1$ is unphysical and hence, the radion potential can only possess a minimum. However, if the bulk scalar is taken to be unstable, then the resulting radion potential can possess either a maximum or a minimum. For the $V^{\prime\prime}=0$ case, the same analysis follows except then, we would have to take the first non-zero term in the Taylor expansion of the bulk potential.
We can thus conclude that if the bulk scalar potential has a global minimum and the corresponding radion potential possesses an appropriate  [^7] Appropriate means (here and further) that the modulus value $r_{c}$ should be positive. extremum, then radion stabilization is guaranteed. However, for bulk scalar potentials with a global maximum and no global minimum, and the corresponding radion potential having an appropriate extremum, we saw that modulus stabilization may or may not be achieved depending on the choice of our parameters for the bulk potential. In this context, if stability (satisfying the BF bound) of the bulk field is further taken into account, then stable fields will always stabilize the radion, and unstable fields may or may not stabilize the radion. An important takeaway is thus the fact that failing to stabilize the radion can only result from an unstable bulk field.
@@
6 General mathematical conditions for the bulk potential to ensure modulus stabilization
@@
In this section, we will frequently use tools from Sec.3. Firstly, note that for a general bulk scalar potential $V(\Phi)$ and defining $z=r_{c}\phi$, the general scalar solution is given by
@@
We write $V^{\prime}(\Phi)|_{\pi r_{c}}$ as $V^{\prime}(\Phi)$ and $V^{\prime}(\Phi_{OR})|_{\pi r_{c}}$ as $V^{\prime}(\Phi_{OR})$ for brevity. Then, plugging these along with the brane potentials defined in Sec.3 into Eq.(16), applying Leibniz’s rule to get the first derivative of the radion potential $V_{GW}(r_{c})$ and setting it to zero (we demand the existence of an extremum at $r_{c}^{*}$) gives us the following relation evaluated at $r_{c}^{*}$
$\begin{split}\frac{\pi}{2}k^{5}\alpha^{2}+\frac{\pi(V^{\prime}(\Phi_{OR}))^{2}}{32k^{2}r_{c}^{*2}}+\frac{\pi k^{3/2}\alpha V^{\prime}(\Phi_{OR})}{4r_{c}^{*}}-\pi k^{5/2}\alpha V^{\prime}(\Phi)-\frac{\pi V^{\prime}(\Phi)V^{\prime}(\Phi_{OR})}{4kr_{c}^{*}}\\ +2\pi\alpha^{2}k^{5}-8\alpha\pi k^{7/2}\Phi_{OR}(\pi r_{c}^{*})+2\alpha k^{5/2}\frac{d\Phi_{OR}(\pi r_{c})}{dr_{c}}|_{r_{c}^{*}}=0\end{split}$ (65)
@@
We now evaluate $V^{\prime\prime}_{GW}(r_{c})$ at $r_{c}^{*}$.
$\begin{split}V^{\prime\prime}_{GW}(r_{c})|_{r_{c}^{*}}=e^{-4kr_{c}^{*}\pi}[\frac{V^{\prime\prime}(\Phi_{OR})}{4r_{c}^{*}}.\frac{d\Phi_{OR}(\pi r_{c})}{dr_{c}}-kV^{\prime\prime}(\Phi)\frac{d\Phi(\pi r_{c})}{dr_{c}}\\ -8k^{2}\frac{d\Phi_{OR}(\pi r_{c})}{dr_{c}}+\frac{2k}{\pi}\frac{d^{2}\Phi_{OR}(\pi r_{c})}{dr_{c}^{2}}-\frac{1}{4r_{c}^{*2}}V^{\prime}(\Phi_{OR})]\end{split}$ (67)
@@
$\begin{split}\frac{2k}{\pi}\frac{d^{2}\Phi_{OR}(\pi r_{c})}{dr_{c}^{2}}|_{r_{c}^{*}}>[8k^{2}+kV^{\prime\prime}(\Phi)|_{r_{c}^{*}}-\frac{V^{\prime\prime}(\Phi_{OR})|_{r_{c}^{*}}}{4r_{c}^{*}}]\frac{d\Phi_{OR}(\pi r_{c})}{dr_{c}}|_{r_{c}^{*}}+\frac{1}{4r_{c}^{*2}}V^{\prime}(\Phi_{OR})|_{r_{c}^{*}}\end{split}$ (68)
@@
$\begin{split}(\frac{2k}{\pi})^{2}\frac{d^{2}\Phi_{OR}(\pi r_{c})}{dr_{c}^{2}}|_{r_{c}^{*}}>[8k^{2}+kV^{\prime\prime}(\Phi)-\frac{V^{\prime\prime}(\Phi_{OR})}{4r_{c}^{*}}]\\ [kV^{\prime}(\Phi)+8k^{2}\Phi_{OR}(\pi r_{c}^{*})-\frac{V^{\prime}(\Phi_{OR})}{4r_{c}^{*}}-\frac{5}{2}\alpha k^{7/2}]+\frac{k}{2\pi r_{c}^{*2}}V^{\prime}(\Phi_{OR})\end{split}$ (69)
@@
where on the right hand side, the “evaluated at $r_{c}^{*}$ is implicit. As one can see, the Eqs.(66) and (69) are completely determined by the bulk scalar potential (note that $\Phi(z)$ depends only on $V^{\prime}(\Phi)$). Thus, these are the two relations that any general bulk potential must satisfy in order for the radion potential to have a stable minimum. It should again be noted that these results hold under the imposed boundary conditions.